CREAM Flight 2008
CSBF Integration Status, July 18, 2008
- Details
Dear colleagues,
Here is a summary of today's activites at CSBF.
- Muon run in flight mode during overnight was running w/o stopping. Data copied to cdps3.
- Nick has looked at the muon data with the well tuned threshold values by Isaac and the TCD signals seem more reasonable even though MXR problem still prevent for the final conclusion.
- Chercam HSK display problem debugging continue. Brian and Jeremy confirmed that the connector was not modified during the test. Next step is to look at the HSK USB output directly.
CSBF Integration Status, July 16, 2008
- Details
Dear colleagues,
Here is a summary of today's activites at CSBF.
- The MXR problem did not disappear during the overnight run. Jim came in the morning and confirmed which cable causes the time out error. The cable lengths between concentrators are checked and sent to Nick. A newly made cable (10 pin flat twisted cable) by Isaac&Nick has been sent via FedEx priority.
- SCD heat pipe fitting was given to Fritz. He has spent time to figure out how to install them. Dave and Fritz talked with Scott on his trip schedule for the installation.
- -X side, sparsification boards stack and master trigger board, has cables cleaned up. So besides SFCs which will be done after the SSDs delivered and integrated, all cablings of instrument are cleaned. I uploaded pictures of cabling on the black board under CSBF integration.
CSBF Integration Status, July 14, 2008
- Details
Dear colleagues,
Here is a summary of today's activities at CSBF.
- Jim re-clamped the flat cable connector on the MXR box. The error message does not show up after that in test mode whole day. But as soon as the muon run in flight mode started in a few minutes, it came back. We will see if this disappears as time goes on as in last two nights.
- Thermal/Vacuum chamber was shown by Joseph. The specification is summarized and uploaded on the black board.
- ZLo puzzle solved by replacing the LVDS cable, which was used from at least last flight. Now the ZLo trigger behaves as we expected.
CSBF Integration Status, July 12, 2008
- Details
Dear colleagues,
Here is a summary of today's activities at CSBF.
- CPT of instrument started. This time, SangEun is the one who excuted all the command and recording.
- CPT of SFC1/2 finished. Some of procedures in the manual need to be corrected and will be done.
- CPT of most of common electronics boards done up to the number 7.n. before the testing of the prescaler.
- Housekeeping board 2ndary boot source did not work during the CPT. Prashant will work on this with Larry's help.
CSBF Integration Status, July 11, 2008
- Details
Dear colleagues,
Here is a summary of today's activities at CSBF.
- Muon run in flight mode during last night stopped as soon as the Chercam LED run started due to the missing delay value table. Fixed.
- Chercam/TCD/CAL groups replied that LPT data look OK so far except the mxr, Chercam housekeeping data, and CAL pedestal noise with HV on.
- Chercam housekeeping display and hk file sent to Laurent for checking and turned out the MUX control cable connector had cable off. Will repair it.



